# Advances in Protein Representation Learning: Methods, Applications, and Future Directions

### v1

@@ Abstract
IR,NMR,UV
1 1. Introduction
Proteins are fundamental biomolecules that are responsible for a wide range of biological processes, including enzymatic catalysis, structural support, signal transduction, and molecular recognition. Their function is dictated by a hierarchical structure comprising four levels: primary (amino acid sequence), secondary (local folding patterns), tertiary (three-dimensional conformation), and quaternary (multi-subunit assembly), as shown in Fig. 1. Understanding these structures is essential for applications in molecular biology, drug discovery, and biotechnology. Understanding protein structure and function is essential to advance molecular biology, drug discovery, and biotechnology. However, the vast diversity of proteins and the complexity of their interactions pose significant challenges to traditional computational and experimental approaches. Protein Representation Learning (PRL) has emerged as a transformative tool for addressing these challenges by encoding protein sequences, structures, and functions into compact, meaningful representations that can be used for various predictive and generative tasks.
Figure 1: The four levels of protein structure, organized by increasing complexity within the polypeptide chain. Primary structure refers to the specific sequence of amino acids. Secondary structure involves local folding patterns, such as $\alpha$-helices and $\beta$-sheets. Tertiary structure represents the overall three-dimensional conformation of a single polypeptide chain. Quaternary structure describes the assembly and interactions of multiple polypeptide chains within a protein complex.
Over the past decade, PRL has witnessed remarkable advancements, fueled by innovations in machine learning and the growing availability of large-scale protein datasets. From feature-based encoding methods to advanced multimodal frameworks that integrate sequence, structure, and functional data, PRL has enabled significant breakthroughs in protein property prediction, structure modeling, and design. These developments have not only advanced fundamental protein science but also opened new avenues for applications in therapeutic development, enzyme engineering, and biomolecular design.
Although numerous surveys have been conducted in the field of PRL, most have been developed from the perspective of biological applications, model architectures, or pretext tasks 1, 2, 3. Although these reviews provide valuable information, they often focus on specialized aspects of PRL, making them less accessible and applicable to researchers outside of AI and computer science. In contrast, our review categorizes PRL research based on the specific modalities utilized, such as sequence, structure, and function, offering a structured framework for understanding how these data types are applied in different approaches. As illustrated in Fig. 2, this organization provides a clear pathway for researchers in various fields to identify PRL methods that align with their data types and research needs. By emphasizing the role of each modality and its relevance in diverse domains, this review bridges the gap between computational advancements and real-world applications, serving as a valuable resource for interdisciplinary researchers seeking to leverage PRL effectively.
In general, the structure of this paper is organized as follows: it begins with Section 1. Introduction, which provides context on PRL and its significance in understanding the structure and function of the protein. Section 2. Feature-Based Approaches explores early methods that encode proteins on the basis of their physical properties. Section 3. Sequence-Based Approaches discusses both non-aligned and aligned sequence methods for capturing protein sequence information. Section 4. Structure-Based Approaches delves into residue-level, atomic-level, protein surface representations, along with symmetry-preserving and equivariant representation learning, emphasizing the critical role of structural data in PRL. Section 5. Multimodal-Based Approaches examines frameworks that integrate sequence, structure, and functional data to generate enriched representations. Section 6. Complex-Based Approaches covers protein-ligand and protein-protein complex representations, focusing on modeling molecular interactions. To facilitate comparison, we provide a structured overview of the strengths and limitations of each approach in Table 1, offering insights to guide the selection of appropriate PRL strategies based on specific biological and computational requirements. Section 7. Databases for Protein Representation Learning introduces key resources for sequences, structures, and functions that support PRL. Section 8. Applications highlights practical implementations of PRL, including protein property prediction, protein structure prediction, protein design and optimization (e.g., ligand-binding proteins, enzymes, and antibodies), and structure-based drug design. Section 9. Future Directions and Open Challenges discusses key challenges such as expanding PRL to DNA / RNA representation learning, improving the scalability and generalization of the model, and improving the explainability, with the aim of inspiring future research and advances in the field. Finally, Section 10. Conclusion provides a summary of the key insights from this review and reflects on the broader impact of PRL, emphasizing its potential for advancing protein science and biomedical applications.
Figure 2: An overview of the key components and themes discussed in this review. The figure highlights the interconnections and relationships between the main topics, providing a comprehensive visual summary of the review’s scope.
2 2. Feature-Based Approaches
Early computational representations of proteins relied on encoding amino acid sequences using physicochemical and biochemical properties 4, 5. This approach assigns numerical attributes—such as charge, hydrophobicity, and molecular size—to amino acids, often sourced from standardized databases like AAIndex 6. These numerical feature vectors enable traditional machine learning models to analyze protein properties based on their physicochemical traits.
Despite their utility, feature-based methods face challenges 7: (i) Selecting relevant properties is non-trivial, as molecular behavior is highly context-dependent and often poorly understood; and (ii) Handcrafted feature engineering may oversimplify protein characteristics, limiting its ability to capture complex sequence-function relationships. As a result, these representations often struggle to generalize across diverse protein-related tasks.
To address these limitations, modern PRL approaches have shifted toward data-driven methods that automatically extract informative representations from raw protein sequences or structures. By leveraging machine learning, these techniques eliminate the need for manual feature selection, improving scalability and predictive performance in applications such as drug discovery, functional annotation, and protein engineering.
3 3. Sequence-Based Approaches
Protein sequences, composed of linear chains of amino acids, serve as the foundation for understanding protein behavior. These sequences can be viewed as a unique “biological language”, where amino acids act as the “words” arranged in a linear “sentence”. This analogy to natural (textual) language has driven the adoption of methodologies from natural language processing (NLP) for PRL, enabling models to capture the implicit “grammar” encoded within protein sequences. Sequence-based approaches focus on extracting meaningful representations directly from these sequences and can be broadly categorized into two main methods: aligned approaches, which utilize evolutionary information from Multiple Sequence Alignments (MSAs), and non-aligned approaches, which process individual sequences independently.
3.1 3.1. Aligned Sequence Approaches
Aligned sequence methods leverage evolutionary relationships by analyzing Multiple Sequence Alignments (MSAs), which align homologous sequences to highlight conserved regions critical for protein function. The long-standing practice in computational biology is to infer structural and functional constraints from evolutionarily related sequences, as conserved regions often correspond to functionally or structurally important sites. Consequently, several protein representation models have been developed to capture co-evolutionary information by taking MSAs as input. These approaches have been particularly effective for protein folding and structure prediction, where evolutionary information plays a critical role 8, 9, 10, 11.
One of the most influential works leveraging MSAs is AlphaFold 8, which marked a significant breakthrough in addressing the protein folding problem. By delivering unprecedented accuracy in predicting protein structures, AlphaFold fundamentally advanced the field and established a foundation for numerous subsequent research endeavors, inspiring innovations in protein structure prediction and related applications.
However, MSA-based methods have limitations. First, they rely on the availability of homologous sequences, which may not exist for every protein 12. Second, some proteins, such as those from orphan families or de novo-designed proteins, lack sufficient evolutionary data, limiting the applicability of these models 12. Additionally, the computational cost of constructing MSAs for large protein families can be prohibitive, leading researchers to explore non-aligned sequence approaches 13.
3.2 3.2. Non-Aligned Sequence Approaches
Non-aligned methods analyze protein sequences independently, without incorporating evolutionary relationships. These approaches are particularly useful for proteins without sufficient homologous sequences and provide an alternative when MSAs are computationally infeasible.
Early non-aligned approaches relied on handcrafted sequence features, such as amino acid composition, k-mer frequency, physicochemical properties, and evolutionary scoring matrices. These features were fed into classical machine learning models, including Support Vector Machines (SVMs) and Random Forests, for protein classification and function prediction 14, 15.
The advent of deep learning led to the adoption of Variational Autoencoders (VAEs), which learn low-dimensional latent representations of protein sequences, enabling efficient exploration of sequence space 16, 17. Other works employed Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks to capture sequential dependencies within protein sequences, improving sequence modeling tasks 18. Additionally, unsupervised Doc2Vec embedding models have been used to learn fixed-length representations of protein sequences by treating them as “documents” composed of “words” 7.
More recently, the field has progressed toward transformer-based protein language models, driven by advancements in large language models (LLMs) from natural language processing. These Protein Language Models (PLMs), such as Evolutionary Scale Modeling (ESM) series 19, 13, 20, ProtTrans 21, ProteinBERT 22, and ProLLaMA 23, have been pre-trained on large-scale protein sequence datasets. Unlike earlier models, PLMs capture contextual relationships between amino acids, enabling the generation of rich embeddings that encode biologically meaningful information. One major advantage of PLMs is that they generalize beyond known protein families, making them effective for de novo protein design and function prediction.
Beyond general-purpose PLMs, specialized models have been developed for specific functional protein families, such as antibody-specific models 24, 25, 26, 27 and enzyme-focused models 28 trained to capture catalytic properties and functional specificity. These specialized PLMs incorporate domain-specific inductive biases, making them more effective for tasks such as antibody affinity maturation, enzyme engineering, and therapeutic protein design.
4 4. Structure-Based Approaches
Structure-based approaches capture the three-dimensional organization of proteins which is crucial for understanding their structural and functional properties. These approaches leverage structural data at various levels of granularity, as illustrated in Fig. 3, including atoms, residues, and surfaces, to provide richer insights into protein stability, function, and interactions. Additionally, they integrate symmetry principles to enhance learning efficiency and improve representation quality.
[…]
4.1 4.1. Residue-Level Representation
@@
Residue-level representation models focus on capturing structural information at the amino acid level. While sequence-based representations are limited to linear order, residue-level approaches consider the spatial arrangement of residues in the folded protein. This is particularly important because residues that are distant in the primary sequence may be brought into close proximity in the three-dimensional structure due to protein folding. By explicitly modeling these spatial relationships, residue-level representations enable a more comprehensive understanding of protein stability, interactions, and functional properties.
@@
Several notable studies have advanced residue-level representations by leveraging graph-based approaches 29, 30, 31, 32, 33, 34, 35, 36, 37, 38. In this representation, proteins are modeled as graphs, where nodes correspond to residues, typically represented by the alpha carbon ($C\alpha$) of each amino acid along with its 3D coordinates. Node features frequently include the amino acid type, physicochemical properties (e.g., charge, hydrophobicity, and polarity), and secondary structure annotations, providing rich information about each residue’s role and environment. The edges of the residue-level graphs vary to reflect different relationships between residues 36:
@@
4.2 4.2. Atomic-Level Representation
Atomic-level representation models capture the fine-grained details of protein structures by encoding the positions and interactions of individual atoms, offering a more detailed view than residue-level approaches. While residue-level representations typically approximate each amino acid with a single point - often the alpha carbon ($C\alpha$) - atomic-level models consider all atoms, including side chains and backbone atoms, preserving essential structural and chemical properties. Side-chain atoms play a fundamental role in molecular recognition, enzymatic activity, and ligand binding, where interactions occur at the atomic scale. By explicitly modeling all atoms, atomic-level representations enhance the ability to predict fine-grained molecular interactions.
Several notable studies have advanced atomic-level representations by leveraging graph-based 34, 39, 40 and point cloud-based approaches 41, 42. In graph-based models, proteins are represented as atomic-level graphs, where nodes correspond to individual atoms, each characterized by features such as atomic type, physicochemical properties, and the types of amino acids that the atom belongs to. Edges capture interactions between atoms, commonly defined based on chemical bonds, non-covalent interactions, or k-nearest neighbors (KNN), which connect each atom to its closest spatial neighbors in 3D space. The KNN-based approach, which is the most commonly used edge construction method in atomic-level graphs, allows models to incorporate both bonded and non-bonded atomic interactions, improving their ability to capture structural and functional relationships. In contrast, point cloud-based representations treat atoms as independent points in three-dimensional space, discarding explicit connectivity but preserving raw spatial distributions. This approach enables models to learn atomic-level geometric features without requiring predefined graph structures, making them particularly effective for capturing fine-grained structural details and local atomic interactions.
4.3 4.3. Protein Surface Representation

### v2

@@ Abstract.
: Corresponding Author
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 24th International Workshop on Data Mining in Bioinformatics; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Representation learning offers a powerful framework for converting high-dimensional, complex data into compact and informative embeddings that retain essential patterns while enabling efficient computation. In the context of protein science, where sequences and structures exhibit immense diversity and complexity, such representations are crucial for supporting both predictive and generative tasks. Proteins are essential biomolecules that drive nearly all cellular processes, from catalyzing biochemical reactions to providing structural integrity and mediating signal transduction. A deep understanding of proteins is critical for numerous scientific and medical applications, including drug discovery, disease mechanism elucidation, and the design of novel enzymes or therapeutics. Central to this understanding is the recognition that the function of a protein is intimately related to its structure, which is organized hierarchically into four levels: primary (amino acid sequence), secondary (local folding motifs), tertiary (three-dimensional conformation) and quaternary (multi-subunit assembly), as illustrated in Fig. 1.
Figure 1. The four levels of protein structure, organized by increasing complexity within the polypeptide chain. Primary structure refers to the specific sequence of amino acids. Secondary structure involves local folding patterns, such as $\alpha$-helices and $\beta$-sheets. Tertiary structure represents the overall three-dimensional conformation of a single polypeptide chain. Quaternary structure describes the assembly and interactions of multiple polypeptide chains within a protein complex.
Despite decades of research, decoding the structure–function relationship remains a profound scientific challenge. The mapping from sequence to structure is highly nonlinear and context-dependent, with even minor sequence variations potentially leading to significant changes in folding and function. Furthermore, proteins operate in dynamic, heterogeneous cellular environments, which adds additional complexity to their accurate modeling. The vast combinatorial space of possible sequences and conformations renders exhaustive experimental characterization infeasible and stretches the limits of traditional computational techniques. In response, Protein Representation Learning (PRL) has emerged as a transformative approach. By encoding protein sequences, structures, and functions into expressive fixed-length vectors, PRL enables scalable and data-driven analysis. Leveraging advances in deep learning and large-scale protein datasets, PRL models have achieved state-of-the-art results in structure prediction, function annotation, and de novo protein design. These developments are reshaping our ability to model, interpret, and engineer proteins with unprecedented precision and scale.
Although numerous surveys have been conducted in the field of PRL, most have been developed from the perspective of biological applications, model architectures, or pretext tasks (wu2022survey, ; xiao2025protein, ; heinzinger2025teaching, ). Although these reviews provide valuable information, they often focus on specialized aspects of PRL, making them less accessible and applicable to researchers outside of AI and computer science. In contrast, our review categorizes PRL research based on the specific modalities utilized, such as sequence, structure, and function, offering a structured framework for understanding how these data types are applied in different approaches. As illustrated in Fig. 2, this organization provides a clear pathway for researchers in various fields to identify PRL methods that align with their data types and research needs. By emphasizing the role of each modality and its relevance in diverse domains, this review bridges the gap between computational advancements and real-world applications, serving as a valuable resource for interdisciplinary researchers seeking to leverage PRL effectively.
In general, the structure of this paper is organized as follows: it begins with Section Introduction, which provides context on PRL and its importance in understanding the structure and function of the protein. Section Feature-Based Approaches explores the early methods that encode proteins on the basis of their physical properties. Section Sequence-Based Approaches discusses both non-aligned and aligned sequence methods for capturing protein sequence information. Section Structure-Based Approaches delves into residue-level, atomic-level protein surface representations, along with learning of symmetry-preserving and equivariant representations, highlighting the critical role of structural data in PRL. Section Multimodal-Based Approaches examines frameworks that integrate sequence, structure, and functional data to generate enriched representations. Section Complex-Based Approaches covers the representations of protein-ligand and protein-protein complexes, focusing on the modeling of molecular interactions. To facilitate comparison, we provide a structured overview of the strengths and limitations of each approach in Table 1, offering insights to guide the selection of appropriate PRL strategies based on specific biological and computational requirements. Section Databases for Protein Representation Learning introduces key resources for sequences, structures, and functions that support PRL. Section Applications highlights practical implementations of PRL, including protein property prediction, protein structure prediction, protein design and optimization (e.g., ligand-binding proteins, enzymes and antibodies), and structure-based drug design. Section Future Directions and Open Challenges discusses key challenges such as expanding PRL to DNA / RNA representation learning, improving the scalability and generalization of the model, and improving the explainability, with the aim of inspiring future research and advances in the field. Finally, Section Conclusion provides a summary of the key insights from this review and reflects on the broader impact of PRL, emphasizing its potential to advance protein science and biomedical applications.
Figure 2. An overview of the key components and themes discussed in this review. The figure highlights the interconnections and relationships between the main topics, providing a comprehensive visual summary of the review’s scope.
2. Feature-Based Approaches
Feature-based protein representations encode sequences in structured numerical vectors by leveraging predefined biochemical, structural, or statistical properties of amino acids. These hand-crafted descriptors have historically played a fundamental role in computational biology, enabling a wide range of machine learning applications, including protein classification, subcellular localization, and prediction of molecular interactions. To facilitate the systematic extraction and application of these descriptors, researchers have developed software frameworks such as iFeature (chen2018ifeature, ) and PyBioMed (dong2018pybiomed, ). These toolkits offer unified implementations of diverse feature encoding schemes, along with integrated support for feature selection, clustering, and dimensionality reduction. Widely adopted descriptors, grouped by the types of information they capture, include:
  • Composition-Based Descriptors: These encodings quantify the overall amino acid content of a sequence without accounting for residue order. Representative examples include Amino Acid Composition (AAC), which computes the normalized frequency of each of the 20 standard amino acids, offering a simple but informative global feature vector. Dipeptide Composition (DPC) extends this idea by measuring the frequencies of all contiguous amino acid pairs (400 possible combinations), capturing short-range sequence patterns that may reflect local structural motifs or functional sites.
  • Sequence-Order Descriptors: These descriptors integrate sequential dependencies between residues to better capture biological relevance. A commonly used example is Pseudo-Amino Acid Composition (PseAAC) (shen2008pseaac, ), which enhances traditional AAC by including correlation factors based on physicochemical properties between residues in various lags. This allows the model to retain information on sequence order and residue interactions, making it particularly effective for tasks such as subcellular localization and functional classification.
  • Evolutionary Descriptors: These features encode residue conservation patterns by aligning a query sequence with homologous sequences in a database. The resulting Position-Specific Scoring Matrix (PSSM) (altschul1997gapped, ) represents the likelihood of observing each amino acid at every position, capturing functionally important and evolutionarily conserved regions. PSSM-derived features are widely used in structure prediction, active site identification, and protein family classification.
  • Physicochemical Descriptors: These are based on experimentally measured or computationally derived amino acid properties to capture biochemical characteristics of the sequence. AAindex (kawashima2000aaindex, ) descriptors use numerical scales, such as hydrophobicity, polarity, isoelectric point and steric parameters, to represent individual residues. These values are then aggregated across the sequence using statistical measures (e.g., mean, standard deviation, autocorrelation). Z-scales (sandberg1998new, ) offer a compact alternative by summarizing multiple physicochemical indices into five orthogonal dimensions using principal component analysis, providing a low-dimensional yet expressive representation.
Although feature-based methods have played an important role in early protein modeling, they have notable limitations (yang2018learned, ). One of the primary challenges lies in selecting the appropriate physicochemical properties to represent amino acids, given the vast number of available descriptors and the fact that the molecular factors underlying protein function are often poorly characterized, task-specific, and highly context-sensitive. This makes it difficult to determine which features are likely to be informative for a given predictive objective. Furthermore, these hand-made representations lack contextual awareness: Each amino acid is encoded independently of its sequence or structural surroundings, making it difficult to capture cooperative effects or long-range dependencies. In contrast, recent deep learning approaches generate rich contextual embeddings that adapt to the position and environment of residues within full sequences or structures. These data-driven models reduce the need for manual feature design and have shown improved generalization across a wide range of protein-related tasks.
3. Sequence-Based Approaches
Protein sequences, composed of linear chains of amino acids, represent the primary structure of proteins and encode the fundamental information that determines their higher-order structures and functions. These sequences can be viewed as a “biological language,” where the order and context of amino acids carry implicit rules and patterns. Inspired by this analogy, sequence-based protein representation learning (PRL) methods adapt techniques from natural language processing to learn the statistical “grammar” embedded within sequences. By training on large-scale sequence data, these models map proteins into latent vector spaces, where the resulting embeddings reflect meaningful biological relationships. Sequence-based approaches can be broadly categorized into aligned methods, which incorporate evolutionary information through Multiple Sequence Alignments (MSAs), and non-aligned methods, which learn directly from individual sequences.
3.1. Aligned Sequence Approaches
Aligned sequence methods leverage evolutionary relationships by analyzing Multiple Sequence Alignments (MSAs), which align amino acid sequences of homologous proteins from different organisms to identify conserved residues. An MSA arranges these sequences into a matrix where each row represents a protein and each column corresponds to a specific residue position, enabling systematic comparison of residue variability and conservation. The core insight behind MSAs is that conserved positions often reflect evolutionary constraints, such as structural stability, enzymatic activity, or interaction specificity, since mutations at these sites are typically deleterious and thus selected against. Consequently, MSAs encode rich contextual information that captures residue co-evolution, functional motifs, and structural dependencies, making them a powerful tool for studying protein behavior.
This evolutionary signal has made MSAs a valuable input for computational models aimed at inferring protein structure and function. Recent work has successfully used MSAs to learn biologically meaningful representations by modeling co-evolutionary patterns and conserved sequence features (rao2021msa, ; jumper2021highly, ; zheng2024improving, ; chen2024msagpt, ). Among these, two of the most influential and foundational are the MSA Transformer and AlphaFold. MSA Transformer (rao2021msa, ) is a representation learning model that processes MSAs using axial attention, a mechanism that applies self-attention along both the rows (sequences) and columns (residue positions) of the alignment. This allows the model to jointly capture patterns of conservation and variation across homologous sequences, resulting in expressive protein embeddings grounded in evolutionary context. AlphaFold (jumper2021highly, ), although not a representation learning model in the strict sense, achieved a major breakthrough in structure prediction by deeply integrating MSA-derived information. Its Evoformer module refines MSA and pairwise residue representations through a series of attention-based communication blocks, enabling the model to infer spatial and functional constraints with remarkable accuracy. These two models are among the earliest and most widely recognized for their effective use of MSA data, and they have laid the foundation for many subsequent advances in protein representation learning and structure prediction.
Despite their effectiveness, MSA-based approaches have notable limitations. Their success depends on the availability of homologous sequences, which may be limited or absent for orphan proteins or de novo-designed sequences (10.1093/bib/bbad217, ). Moreover, constructing high-quality MSAs can be computationally expensive, especially at scale. These constraints have prompted increasing interest in non-aligned sequence methods that learn directly from individual sequences without relying on explicit evolutionary data (lin2023evolutionary, ).
3.2. Non-Aligned Sequence Approaches
Non-aligned sequence methods learn directly from individual protein sequences without relying on Multiple Sequence Alignments (MSAs) or explicit evolutionary context. These approaches are particularly valuable for proteins lacking sufficient homologs and are well suited for large-scale modeling, where the construction of MSAs can be computationally infeasible. In contrast to MSA-based methods, which explicitly leverage homologous sequences to capture evolutionary conservation, non-aligned models learn implicit biological constraints from the statistical co-occurrence of residues across large and diverse sequence corpora. This allows them to generalize more flexibly to proteins with limited or no evolutionary annotations, including those from underrepresented families or synthetic origins.
Deep learning has enabled a shift toward data-driven methods that learn representations from raw sequences. Early neural models such as Variational Autoencoders (VAEs) (sinai2017variational, ; ding2019deciphering, ) captured the generative structure of protein space by mapping sequences to low-dimensional latent vectors. Recurrent architectures such as RNN and LSTM further advanced this line of work by modeling long-range sequence dependencies (almagro2020language, ). A notable milestone was UniRep (alley2019unified, ), which introduced a self-supervised approach using mLSTM trained on millions of sequences. UniRep demonstrated that fixed-length embeddings could predict a range of biochemical properties, laying the foundation for the era of protein language models (PLMs).
More recently, transformer-based Protein Language Models (PLMs) have become the dominant paradigm, drawing inspiration from large language models in natural language processing. Models such as the Evolutionary Scale Modeling (ESM) series (rives2021biological, ; lin2023evolutionary, ; hayes2024simulating, ), ProtTrans (elnaggar2021prottrans, ), ProteinBERT (brandes2022proteinbert, ), and ProLLaMA (lv2024prollama, ) are pre-trained on massive protein sequence datasets using masked language modeling objectives. During training, PLMs learn contextual dependencies by predicting masked amino acids, resulting in internal embeddings that encode rich, task-agnostic information about sequence structure and function. These representations form a learned latent space, where similar proteins are embedded closer together, enabling knowledge transfer across diverse sequence types. In addition to general-purpose PLMs, specialized models have been developed to target particular protein classes. For example, antibody-specific PLMs (olsen2022ablang, ; gao2023pre, ; shuai2023iglm, ; kenlay2024large, ) and enzyme-focused models (munsamy2022zymctrl, ) incorporate domain-specific inductive biases to improve performance in tasks such as affinity maturation and prediction of catalytic function.
4. Structure-Based Approaches
Structure-based approaches capture the three-dimensional organization of proteins, which is crucial to understanding their structural and functional properties. These approaches leverage structural data at various levels of granularity, as illustrated in Fig. 3, including atoms, residues, and surfaces, to provide richer insight into protein stability, function, and interactions. In addition, they integrate symmetry principles to improve learning efficiency and improve representation quality.
@@ 4. Structure-Based Approaches
4.1. Residue-Level Representation
@@
Residue-level representation models focus on capturing structural information at the amino acid level. Although sequence-based representations are limited to linear order, residue-level approaches consider the spatial arrangement of residues in the folded protein. This is particularly important because residues that are distant in the primary sequence may be brought into close proximity in the three-dimensional structure due to protein folding. By explicitly modeling these spatial relationships, residue-level representations enable a more comprehensive understanding of protein stability, interactions, and functional properties.
@@
Several notable studies have advanced residue-level representations by leveraging graph-based approaches (hermosilla2022contrastive, ; xia2022fast, ; zhou2022lightweight, ; le2022representation, ; guo2022self, ; wang2023learning, ; chen20233d, ; zhang2023protein, ; Gao2023, ; voitsitskyi20233dprotdta, ). In this representation, proteins are modeled as graphs, where the nodes correspond to residues, typically represented by the alpha carbon ($C\alpha$) of each amino acid along with its 3D coordinates. Node features frequently include amino acid type, physicochemical properties (e.g. charge, hydrophobicity, and polarity) and secondary structure annotations, providing rich information about each residue’s role and environment. The edges of the residue-level graphs vary to reflect different relationships between residues (zhang2023protein, ):
@@
Beyond graph-based approaches, recent work has explored discretized residue-level representations for scalable structural modeling and comparison (van2024fast, ; zhang2024balancing, ; gaujac2024learning, ). These methods encode local or global 3D structural environments into a finite set of tokens, effectively transforming protein structures into symbolic sequences. Many of these representations are learned using vector-quantized variational autoencoders (VQ-VAEs), which cluster local residue geometries into discrete structural alphabets. A typical and widely adopted example is Foldseek (van2024fast, ), which constructs such an alphabet by assigning each residue a token that summarizes its spatial context relative to neighboring residues. This enables sequence-style processing with efficient alignment while retaining biologically relevant structural information. Unlike models that explicitly define spatial graphs, discretized representations offer compact and alignment-friendly abstractions suitable for large-scale structure analysis and modeling.
4.2. Atomic-Level Representation
Atomic-level representation models capture the fine-grained structural and chemical characteristics of proteins by explicitly modeling individual atoms and their spatial arrangements. This level of granularity preserves critical details necessary to understand molecular interactions, since it includes all atoms within a protein, both backbone and side chains, rather than approximating each residue with a single representative point, such as alpha carbon ($C\alpha$). By incorporating atom-specific features and spatial relationships, atomic-level approaches provide a richer and more precise encoding of the protein structure. Several modeling paradigms have been proposed at this level, including graph-based representations, point cloud methods, and voxel-based 3D convolutional neural networks (3D CNNs), each offering unique inductive biases and trade-offs in terms of expressiveness, efficiency, and spatial resolution.
Graph-based methods represent protein structures as atom-level graphs, where nodes correspond to individual atoms annotated with features such as element type, partial charge, hybridization state, or residue identity (wang2023learning, ; wu2023atomic, ; jiale2024pretraining, ). Edges define pairwise relationships between atoms, typically based on chemical bonds, spatial proximity, or non-covalent interactions. A common strategy is to use $k$-nearest neighbor (KNN) connections to flexibly capture both bonded and non-bonded atomic interactions. These graphs are well-suited for message-passing neural networks and geometric learning architectures that propagate information over the molecular graph to model structural dependencies.
Point cloud-based methods treat atoms as an unordered set of points in three-dimensional space, each point associated with features such as atom type or partial charge (wang2022point, ; nguyen2024multimodal, ). These approaches forego explicit graph connectivity, relying instead on spatial convolutions or attention mechanisms to learn directly from the geometric distribution of atoms. This makes point cloud representations efficient and adaptable, especially for modeling flexible or irregular structures, where fixed connectivity may be limiting. The raw spatial encoding also allows these models to generalize across diverse molecular conformations.
Voxel-based 3D convolutional neural networks (3D CNNs) provide a dense grid-based representation of atomic environments (torng2019high, ; torng20173d, ). In this framework, atoms are projected into a fixed-resolution three-dimensional lattice, where each voxel encodes the presence and chemical identity of atoms within a defined spatial window. These volumetric grids are processed using 3D convolutional layers that capture local geometric and chemical patterns. Unlike graph- and point-cloud models, 3D CNNs operate in structured input spaces, leveraging established computer vision techniques. While computationally more demanding due to their dense spatial encoding, 3D CNNs offer high spatial fidelity and do not require predefined topological constraints.
4.3. Protein Surface Representation